Submitted to the 1. Conference on Designing a L^{A}TEX2e Class for Typesetting ACM Papers, Hawaii 2000

ACM SIG Proceedings Format

1st Author

First author’s affiliation 1st line of address 2st line of address

Telephone number, incl. country code

1st author’s email address

2nd Author

First author’s affiliation 1st line of address 2st line of address

Telephone number, incl. country code

2nd author’s email address

3rd Author

First author’s affiliation 1st line of address 2st line of address Telephone number, incl. country code

3rd author’s email address

ABSTRACT

In this paper, we describe the formatting guidelines for ACM SIG Proceedings.

KEYWORDS

Guides, instructions, authors kit, conference publications.

1 INTRODUCTION

The proceedings are the records of the conference. ACM hopes to give these conference by-products a single, highquality appearance. To do this, we ask that authors follow some simple guidelines. In essence, we ask you to make your paper look exactly like this document. The easiest way to do this is simply to down-load a template from [1], and replace the content with your own material.

2 PAGE SIZE

All material on each page should fit within a rectangle of 18 x 23.5 cm (6” x 9.25”), centered on the page, beginning 2.54 cm (1”) from the top of the page and ending with 2.54 cm (1”) from the bottom. The right and left margins should be 1.9 cm (.75”). The text should be in two 8.45 cm (3.33”) columns with a .83 cm (.33”) gutter.

3 TYPESET TEXT

Prepare your submissions on a typesetter or word processor.

LEAVE BLANK THE LAST 3.81 cm (1.5”) OF THE LEFT COLUMN ON THE FIRST PAGE FOR THE COPY-RIGHT NOTICE

3.1 Normal or Body Text

Please use a 9-point Times Roman font, or other Roman font with serifs, as close as possible in appearance to Times Roman in which these guidelines have been set. The goal is to have a 9-point text, as you see here. Please use sansserif or non-proportional fonts only for special purposes, such as distinguishing source code text. The Press 9-point font available to users of Script is a good substitute for Times Roman. If Times Roman is not available, try the font named Computer Modern Roman. On a Macintosh, use the font named Times. Right margins should be justified, not ragged.

3.2 Title and Authors

The title (Helvetica 18-point bold), authors’ names (Helvetica 12-point) and affiliations (Helvetica 10-point) run across the full width of the page - one column wide. We also recommend phone number (Helvetica 10-point) and email address (Helvetica 12-point). See the top of this page for three addresses. If only one address is needed, center all address text. For two addresses, use two centered tabs, and so on. For more than three authors, you may have to improvise^{1}

3.3 First Page Copyright Notice

Leave 3.81 cm (1.5”) of blank space at the bottom of the left column of the first page for the copyright notice.

3.4 Subsequent Pages

For pages other than the first page, start at the top of the page, and continue in double-column format. The two columns on the last page should be of equal length.

3.5 References and Citations

Footnotes should be Times New Roman 9-point.

Use the standard Communications of the ACM format for references - that is, a numbered list at the end of the article, ordered alphabetically by first author, and referenced by numbers in brackets [2]. See the examples of citations at

1If necessary, you may place some address information in a foortone, or in a named section at the end of your paper. the end of this document [3, 4, 5]. Within this template file, use the style named references for the text of your citation.

References should be published materials accessible to the public. Internal technical reports may be cited only if they are easily accessible (i.e. you can give the address to obtain the report within your citation) and may be obtained by any reader. Proprietary information may not be cited. Private communications should be acknowledged, not referenced (e.g., “[Robertson, personal communication]”).

3.6 Page Numbering, Headers and Footers

Do not include headers, footers or page numbers in your submission. These will be added when the publications are assembled.

4 SECTIONS

The heading of a section should be in Times New Roman 12point bold in all-capitals flush left. Sections and subsequent subsections should be numbered and flush left.

4.1 Subsections

The heading of subsections should be in Times New Roman 12-point bold with only the initial letters capitalized. (Note: For subsections and subsubsections, a word like the or a is not capitalized unless it is the first word of the header.)

4.1.1 Subsubsections

The heading for subsubsections should be in Times New Roman 11-point italic with initial letters capitalized.

4.1.1.1 Subsubsubsection

The heading for subsubsubsections should be in Times New Roman 11-point italic with initial letters capitalized.

4.1.1.1.1 Subsubsubsubsection

The heading for subsubsubsubsections should be in Times New Roman 11-point italic with initial letters capitalized.

5 FIGURES/CAPTIONS

Place Tables/Figures/Images in text as close to the reference as possible. It may extend across both columns to a maximum width of 17.78 cm (7”).

Captions should be Times New Roman 9-point bold. They should be numbered (e.g., “Table 1” or “Figure 2”) and be centered beneath each table, figure or image.

6 ACKNOWLEDGMENTS

Our thanks to ACM SIGCHI for allowing us to modify templates they had developed.

REFERENCES

[1] ACM. ACM SIG PROCEEDINGS template. Technical report, ACM, http://www.acm.org/sigs/ pubs/proceed/template.html, 1999.

[2] R.E Anderson. Social impacts of computing: Codes of professional ethics. Social Science Computing Review, 10(4):453–469, 1992.

[3] Sue Conger and Karen D. Loch. Ethics and computer use. Communications of the ACM, 38(12):30–32, December 1995.

[4] W.E. Mackay. Ethics, lies and videotape... In Proceedings of CHI ’95 (Denver CO), pages 138–145. ACM Press, May 1995.

[5] M. Schwartz and Task Force on Bias-Free Language. Guidelines for bias-free writing. Technical report, Indiana University, 1995.

Columns on Last Page Should Be Made Equal Length

1st Author, 31. December 1999  2  CONF-2000